Please help me solve: slope -2/5 passing through the x coordinate 10/4
I am trying to help my Godson but I need to understand how to tell hime how to solve it. Thanks

Answer by Alan3354(69443)

Please help me solve: slope -2/5 passing through the x coordinate 10/4
-------------------
I think you mean (10,4)
----------
Use y = mx + b and the point to find b
4 = (-2/5)*10 + b
4 = -4 + b
b = 8
------
Then it's y = mx + b
y = (-2/5)x + 8